# ForceGrip: Data-Free Curriculum Learning for Realistic Grip Force Control in VR Hand Manipulation

### v1

@@ 4.2.3 Phase 3: Dynamic Stabilization
5 Curriculum and Ablation Study
@@
5.1 Test Conditions
Recall that our curriculumn comprises three sequential phases: Finger Positioning (P1), Intention Adaptation (P2), and Dynamic Stabilization (P3). This aims to decompose randomized factors—such as varying trigger inputs and wrist motions—into progressively challenging phases, thereby improving learning stability in physically simulated VR environments.
@@ 5.1 Test Conditions
Both ablation conditions use the C1 curriculum with the same epoch-based transitions to ensure fair comparisons. In preliminary analyses, we confirmed that these phase-transition epochs did not adversely affect convergence when applied consistently across all setups. Note that, for EL $\times$, voxel-based encoders for $v_{gvs}$ and $v_{lvs}$ were retained to ensure the model’s trainability, given the large feature sizes of these components.
5.2 Evaluation Results
To investigate the training progression, we measured reward in every epoch for each curriculum and ablation setting. Figure 3 illustrates reward curves presenting how the agent’s performance evolves over time. Curricula that do not begin with P1 (C4, C5, and C6) exhibit noticeably lower rewards throughout training. This outcome emphasizes learning stable finger contact before introducing additional complexities, such as random wrist movement or realistic trigger inputs. Among the curricula, our proposed approach (C1) achieves the highest overall reward.
Figure 3: Reward curves for the different curriculum designs (C1C6) and ablation experiments (PR $\times$, EL $\times$). The vertical lines at epochs 1200 and 1600 mark the transitions between Phases 1, 2, and 3. Note that in the PR $\times$ condition, the reward values are doubled because only a single reward term remains once the proximity reward is removed.
For a more in-depth analysis, we collected data from 100,000 episodes and measured three key metrics: episode success ratio (ESR), average proximity reward (PR), and average force reward (FR). Table 2 presents the quantitative results. While C4, C5, and C6 maintain relatively acceptable ESR and PR scores, their lower FR scores reveal difficulties in managing force control when advanced randomization (e.g., full wrist movement or dynamic triggers) is introduced early. By contrast, C2 and C3—which incorporate key phases in a more structured sequence—yield moderate ESR but fall below C1 in overall stability. Removing the proximity reward (PR $\times$) significantly reduces performance, particularly in ESR and PR, but also affects FR. Although eliminating encoder layers (EL $\times$) achieves a higher ESR and PR, it simultaneously lowers the FR score. This outcome suggests that the network, lacking encoder-based feature extraction, struggles to incorporate critical low-dimensional inputs (e.g., $u\in\mathbb{R}^{6}$), causing the agent to adopt an overly forceful grip rather than refining grip force precision.
Table 2: Quantitative evaluation results. The highest score in each row is bold and underlined, and the second-highest is underlined only.
@@ 5.2 Evaluation Results
Overall, these findings confirm that a carefully structured curriculum—initiating with stable finger contact, followed by progressive introduction of force variation and wrist movement—affords both higher success rates and more realistic grip forces. Furthermore, ablation results emphasize the importance of our design choices, such as the proximity reward and encoder layers, in achieving robust and precise grip force control.
6 User Study
User experience is crucial when assessing the effectiveness of VR interaction techniques, especially those involving detailed hand manipulation. To present the superiority of our method, ForceGrip (FG), we conducted a comparative user study against two state-of-the-art VR controller interface methods: Attachment (AT) Oprea et al. (2019) and VR-HandNet (VH)  Han et al. (2023).
6.1 Evaluation Metrics
To evaluate visual realism, we adopted the Realism Questionnaire (RQ) proposed by Oprea et al. (2019) containing 14 items across three categories: motor control, finger movement realism, and interaction realism. The individual scores are then aggregated into category-specific and overall indices, providing insights into perceptual realism from a purely visual standpoint.
To evaluate physical user experience, we devised a Force Realism Questionnaire (FRQ). Inspired by Gonzalez-Franco and Peck (2018) and tailored to physically simulated hand interactions, the FRQ retains the RQ’s three-category framework but focuses on evaluating grip force accuracy and user intention. Detailed FRQ items and scoring procedures are provided in the supplementary material.
To ensure that user comfort did not skew any evaluations, we employed the Simulator Sickness Questionnaire (SSQ) Kennedy et al. (1993) at the beginning and conclusion of each condition.
We applied the Friedman test to detect statistically significant differences, as the data did not meet the normality assumptions required by parametric tests. For post-hoc analysis, we conducted Wilcoxon signed-rank tests.
6.2 Participants and Apparatus
@@
We conducted the experiment using an Oculus Quest 2 headset and a stock controller, powered by a desktop featuring an NVIDIA RTX 3080 GPU and an AMD Ryzen 5 5600G CPU. This setup provided consistent performance across all trials, minimizing hardware-induced variability.
6.3 E1: Pick-and-place — Method
Figure 4: Pick-and-place task. (a) Randomly generated objects with color label pairs (R, G, or B). (b) Procedure: pick up the object, rotate and inspect its color label, and place it into the corresponding color box. (c) 15 candidate objects are displayed at three scales: original, isotropic, and anisotropic.
@@ 6.3 E1: Pick-and-place — Method
6.4 E1: Pick-and-Place — Results
Figure 5: RQ scores for the Pick-and-Place task, using a 7-point Likert scale ranging from -3 (strongly disagree) to +3 (strongly agree). Brackets indicate statistical significances (* : $p$ $<$ .05, **∗ ∗ : $p$ $<$ .01).
@@ 6.4 E1: Pick-and-Place — Results
Participants provided qualitative observations that illuminated each technique’s strengths and limitations. For FG, participants praised it’s ability to “stably and realistically grasp small objects” (P3, P7, P12) and “naturally adapt finger motions to varied shapes” (P6, P9, P11, P13, P14, P19, P20). Some, however, experienced occasional “object slipping” (P3, P6, P10), likely arising from the rigid-body collision implementation of the capsule colliders at the fingertips.
@@
Overall, this experiment indicates that ForceGrip provides superior visual realism for typical object manipulation tasks.
6.5 E2: Can squeeze — Method
Figure 6: Can squeeze task. (a) Generated can object with a random target level. (b) Procedure: pick up the can, squeeze it via trigger control, and place it into the container. (c) Can deformation levels (L0: no deformation - L10: maximum deformation) visualized with mesh representations.
@@ 6.5 E2: Can squeeze — Method
Participants were presented with a virtual can-shaped object modeled after standard beverage cans. They were then asked to deform it to a target level ranging from 0 (no deformation) to 10 (maximum deformation). Each integer increment (from 0-10) corresponded to a 1 kgf increase in force. We adopted this discrete, 11-level scale to balance force granularity with the limited sensitivity of typical VR controller triggers. In preliminary testing, subdividing the force range more finely led to confusion and reduced accuracy, as participants struggled to modulate shallow trigger inputs with greater precision.
At the start of each trial, a can was introduced in front of the user while a random target deformation level appeared on the display. The participant then adjusted the VR controller’s trigger to deform the can until it closely matched the target level, and subsequently placed the deformed can into a designated container. During the interaction, both visual and auditory feedback—implemented through mesh updates and sound effects—indicated the extent of can deformation.
@@
6.6 E2: Can squeeze — Results
Figure 7: FRQ scores for the Can Squeeze task, using a 7-point Likert scale ranging from -3 (strongly disagree) to +3 (strongly agree). Brackets indicate statistical significances (* : $p$ $<$ .05, **∗ ∗ : $p$ $<$ .01).
As shown in Figure 7, there exist significant differences between FG and VH across all metrics. Detailed statistical comparison results are provided in the supplementary materials. There was no statistically significant change in SSQ scores before and after the experiment.
@@ 6.6 E2: Can squeeze — Results
Figure 8: Correlation between trigger input ($x$) and grip force ($y$) for FG and VH, with the reference line ($y=10x$). Outliers (top 5%) were excluded, and the remaining data was duplicated and resampled to 1,000 points, ensuring an even distribution of trigger values. For FG, points below the reference line mostly occurred when the object slipped from the agent’s grasp.
@@
To visualize how closely each system reflects user intentions, we plotted the trigger input against the measured grip force (Figure 8). FG exhibited a clear linear correlation ($r=0.884$). In contrast, VH showed negligible correlation ($r=0.073$), consistently producing excessive grip forces irrespective of trigger amplitude. This result contradicts the 1.2 kgf maximum force claimed by Han et al. (2023) and supports reports that VH appears to lack grip force control.
7 Conclusion and Future Improvements
@@
We have introduced ForceGrip, a novel approach for delivering realistic hand manipulation in VR controller interfaces through faithful grip force control. Unlike many prior systems that rely on narrowly scoped reference datasets, ForceGrip leverages automatically generated training scenarios, thereby improving adaptability and avoiding biases intrinsic to motion-capture data. Our curriculum learning framework further addresses the challenges posed by complex, dynamic interactions, enabling robust convergence and state-of-the-art performance across diverse virtual manipulation tasks.
One significant limitation stems from using standard VR controllers, whose trigger buttons lack fine-grained control and haptic feedback. Future work may explore advanced sensors—such as electromyography, elastic potentiometers, or pressure sensors—augmented by force-feedback mechanisms to offer richer feedback loops and more intuitive control. Beyond hardware enhancements, improving collision and friction modeling through soft-body simulations could produce more stable and lifelike contact dynamics, particularly for heavier or deformable objects.
Appendix A PPO hyperparameters
The parameters for proximal policy optimization (PPO) are shown in Table 3.
Table 3: PPO hyperparameters used in the experiment
Parameter Value Parameter Value
Actor learning rate 0.0001 PPO epochs 16
Critic learning rate 0.001 Episode batch size 3000
Discount factor 0.98 Mini batch size 3000
GAE parameter 0.95
Clipping parameter ($\epsilon$) 0.2
Entropy coefficient 0.01
Appendix B Detailed Network Architecture
Detailed network architecture is illustrated in Figure 9.
Figure 9: Detailed network architecture. FC(nn) represents a fully connected layer with n nodes, and Drop(0.1) denotes a dropout layer with a rate of 0.1.
Appendix C Trigger flow generation
The algorithms used for various trigger flow generation are described in algorithm 1.
Input: Time Range $t_{0}$ to $t_{T}$, Frequency Range $[f_{\text{min}},f_{\text{max}}]$, Amplitude Range $[A_{\text{min}},A_{\text{max}}]$, Offset Range $[O_{\text{min}},O_{\text{max}}]$, Noise Parameters $\sigma$ (std. deviation), clip (max noise magnitude)
Output: Noisy trigger sequence $\{\text{u}_{t}\}_{t=t_{0}}^{t_{T}}$
Step 1: Generate Random Parameters;
$f\leftarrow\text{RandomUniform}(f_{\text{min}},f_{\text{max}})$ ;
  /* Frequency */
$A\leftarrow\text{RandomUniform}(A_{\text{min}},A_{\text{max}})$ ;
  /* Amplitude */
$O\leftarrow\text{RandomUniform}(O_{\text{min}},O_{\text{max}})$ ;
  /* Offset */
$\phi\leftarrow\text{RandomUniform}(0,2\pi)$ ;
  /* Initial Phase */
Step 2: Compute Trigger for Each $t$;
for $t\leftarrow t_{0}$ to $t_{T}$ do
[…]
       $\text{noise}\leftarrow\text{Clamp}(\sigma\cdot\text{RandomGaussian}(),-\text{clip},\text{clip})$;
[…]
end for
return $\{\text{trigger}_{t}\}_{t=t_{0}}^{t_{T}}$;
Algorithm 1 Generate Noisy Trigger Flow
The parameters for noisy trigger flow generation are shown in Table 4.
Table 4: Parameters for Noisy Trigger Flow Generation
     Description / Range      Value
     Frequency range ($[f_{\min},f_{\max}]$)      [0.833, 0.5]
     Amplitude range ($[A_{\min},A_{\max}]$)      [0.0, 2.0]
     Offset range ($[O_{\min},O_{\max}]$)      [-1.0, 1.0]
     Noise standard deviation ($\sigma$)      0.05
     Maximum noise magnitude (clip)      0.1
The examples of the generated trigger flow can be found in Figure 10.
Figure 10: Generated trigger flow examples illustrate the diverse dynamics of noisy synthetic inputs, combining sinusoidal components and Gaussian noise. Each plot represents flow patterns over time (3 s), showcasing behaviors such as gradual activation, periodic oscillations, transient bursts, and steady states, reflecting the natural variability of user behavior.
Appendix D Force Realism Questionnaire
The list of questions in the Force Realism Questionnaire (FRQ) is shown in Table 5. Each item was evaluated using a 7-point Likert scale anchored at ”strongly disagree (-3)” and ”strongly agree (+3).”
Table 5: Force realism questionnaire
ID Question
Aspect 1: Motor Control
Q1. I felt like I could control the grip force as if it were my own hand’s grip force
Q2. The adjustment of the grip force was caused by my controls
Q3. I felt as if the grip force of the virtual hand was influencing my own control
Q4. I felt as if the virtual hand applied grip force by itself
Aspect 2: Grip Force Realism
Q5. It seemed that the grip force change was smooth and plausible
Q6. I felt the grip force was applied and released in a natural way
Q7. Grip forces reacted adequately to my intentions
Aspect 3: Interaction Realism
Q8. I felt like I could adjust the grip force to the object wherever I wanted to
Q9. It seemed as if the virtual fingers were mine when applying grip forces to an object
Q10. I felt that adjusting the grip force was clumsy and hard to achieve
Q11. It seemed as if the grip force was guided and unnatural
Q12. I felt that the grip force was visually correct and natural
Q13. I felt that the grip force was physically correct and natural
Q14. It seemed that the grip force was adapting properly to different geometries
The FRQ consists of three categories, calculated as follows:
[…]
The scoring methodology for the FRQ was adapted from the realism questionnaire (RQ) in Oprea et al. [2019], with the same calculations applied to the RQ’s items as well.
Appendix E Detailed statistical comparison results
E.1 E1: Pick-and-place result
For the pick-and-place experiment, three methods were evaluated: ForceGrip (FG), Attachment (AT) Oprea et al. [2019] and VR-HandNet (VH)  Han et al. [2023].
For the RQ total score, FG scored $1.417\pm 0.862$, VH scored $-0.206\pm 1.497$, and AT scored $0.460\pm 1.050$. A Friedman test revealed a highly significant difference (p-value $=0.0002$, p $<0.01$). For the RQ motor control score, FG had a score of $1.300\pm 1.002$, VH scored $0.325\pm 1.428$, and AT scored $0.588\pm 0.792$, with the Friedman test showing a significant difference (p-value $=0.0080$, p $<0.01$). Regarding the RQ finger movement realism score, FG scored $1.567\pm 1.170$, VH scored $-0.433\pm 1.758$, and AT scored $0.467\pm 1.322$, with a highly significant difference indicated by a Friedman test (p-value $=0.0001$, p $<0.01$). For the RQ interaction realism score, FG scored $1.400\pm 1.024$, VH scored $-0.357\pm 1.691$, and AT scored $0.393\pm 1.268$, with the Friedman test also showing a highly significant difference (p-value $=0.0001$, p $<0.01$).
Pairwise comparisons revealed that for the total score, FG scored significantly higher than VH (p $=0.0001$, p $<0.01$) and AT (p $=0.0027$, p $<0.01$), while the difference between VH and AT was not significant (p $=0.1650$). For the motor control score, FG scored significantly higher than VH (p $=0.0077$, p $<0.01$) and AT (p $=0.0216$, p $<0.05$), while the difference between VH and AT was not significant (p $=0.8955$). In the finger movement realism score, FG was significantly higher than VH (p $=0.0006$, p $<0.01$) and AT (p $=0.0017$, p $<0.01$), with no significant difference between VH and AT (p $=0.0625$). Finally, for the interaction realism score, FG scored significantly higher than VH (p $=0.0004$, p $<0.01$) and AT (p $=0.0032$, p $<0.01$), while VH and AT did not differ significantly (p $=0.0897$).
E.2 E2: Can squeeze result
For the can-squeeze experiment, two methods were evaluated: ForceGrip (FG) and VR-HandNet (VH) Han et al. [2023].
For the FRQ total score, FG scored $1.124\pm 0.838$, while VH scored $-1.357\pm 0.816$, with a highly significant difference (p-value $=0.0000$, p $<0.01$). Regarding the FRQ motor control score, FG achieved $0.900\pm 1.032$ compared to VH’s $-0.613\pm 1.014$ (p-value $=0.0002$, p $<0.01$). For the FRQ grip force realism score, FG scored $1.267\pm 0.981$, surpassing VH’s $-1.900\pm 1.221$ (p-value $=0.0001$, p $<0.01$). Similarly, for the FRQ interaction realism score, FG’s $1.164\pm 1.023$ outperformed VH’s $-1.457\pm 0.941$, demonstrating a highly significant difference (p-value $=0.0000$, p $<0.01$).

### v3

@@ 4.2.3. Phase 3: Dynamic Stabilization
5. Curriculum and Ablation Study
@@
5.1. Test Conditions
Recall that our curriculumn comprises three sequential phases: Finger Positioning (P1), Intention Adaptation (P2), and Dynamic Stabilization (P3). This aims to decompose randomized factors—such as varying trigger inputs and wrist motions—into progressively challenging phases.
@@ 5.1. Test Conditions
Both ablation conditions use the C1 curriculum with identical phase-transition epochs to ensure fair comparison. Preliminary analysis confirmed these transitions did not hinder convergence. For EL $\times$, voxel-based encoders for $v_{gvs}$ and $v_{lvs}$ were retained to maintain trainability due to their large feature sizes.
Figure 4. Reward curves for the different curriculum designs (C1C6) and ablation experiments (PR $\times$, EL $\times$). The vertical lines at epochs 1200 and 1600 mark the transitions between Phases 1, 2, and 3. Note that in the PR $\times$, the curve is doubled to match scale, as only a single reward term remains.
[…]
Line graph comparing training rewards over epochs for six curriculum learning strategies and two ablation conditions. The x-axis shows training epochs, and the y-axis shows normalized reward values. Vertical lines indicate transitions between training phases. ForceGrip with standard curriculum shows the highest convergence.
5.2. Evaluation Results
To investigate the training progression, we measured reward in every epoch for each curriculum and ablation setting (Figure 4). Curricula that do not begin with P1 (C4, C5, and C6) exhibit noticeably lower rewards throughout training. This outcome emphasizes learning stable finger contact before introducing additional complexities. Among the curricula, our proposed approach (C1) achieves the highest overall reward.
Table 2. Quantitative evaluation results. The highest score in each row is bold and underlined, and the second-highest is underlined only.
@@ 5.2. Evaluation Results
[…]
A table showing quantitative evaluation results across six curriculum variants and two ablation studies. Metrics include Episode Success Ratio (ESR), Proximity Reward (PR), and Force Reward (FR). Bold and underlined values indicate the best performance, and underlined values indicate the second best. C1 performs best in force control.
For a more in-depth analysis, we collected data from 100,000 episodes and measured three key metrics (Table 2): episode success ratio (ESR), average proximity reward (PR), and average force reward (FR). While C4, C5, and C6 maintain relatively acceptable ESR and PR scores, their lower FR scores reveal difficulties in managing force control when advanced randomization (e.g., full wrist movement or dynamic triggers) is introduced early. By contrast, C2 and C3—which incorporate key phases in a more structured sequence—yield moderate ESR but fall below C1 in overall stability. Removing the proximity reward (PR $\times$) significantly reduces performance, particularly in ESR and PR, but also affects FR. Although eliminating encoder layers (EL $\times$) achieves a higher ESR and PR, it simultaneously lowers the FR score. This outcome suggests that the network, lacking encoder-based feature extraction, struggles to incorporate critical low-dimensional inputs (e.g., $u\in\mathbb{R}^{6}$), causing the agent to adopt an overly forceful grip rather than refining grip force precision.
Overall, the results confirm that our structured curriculum improves both success rates and grip realism. Ablation studies highlight the impact of design choices such as the proximity reward and encoder layers on precise grip force control.
6. User Study
To present the superiority of our method, ForceGrip (FG), we conducted a comparative user study against two state-of-the-art VR controller interface methods: Attachment (AT) (Oprea et al., 2019) and VR-HandNet (VH)  (Han et al., 2023).
6.1. Evaluation Metrics
To evaluate visual realism, we adopted the Realism Questionnaire (RQ) proposed by (Oprea et al., 2019) containing 14 items across three categories: motor control, finger movement realism, and interaction realism. The individual scores are then aggregated into category-specific and overall indices, providing insights into perceptual realism from a purely visual standpoint.
To evaluate physical user experience, we devised a Force Realism Questionnaire (FRQ). Inspired by (Gonzalez-Franco and Peck, 2018) and tailored to physically simulated hand interactions, the FRQ retains the RQ’s three-category framework but focuses on evaluating grip force accuracy and user intention. Detailed FRQ items and scoring procedures are provided in the supplementary material.
To ensure that user comfort did not skew any evaluations, we employed the Simulator Sickness Questionnaire (SSQ) (Kennedy et al., 1993) at the beginning and conclusion of each condition.
We applied the Friedman test to detect statistically significant differences, as the data did not meet the normality assumptions required by parametric tests. For post-hoc analysis, we conducted Wilcoxon signed-rank tests. Detailed statistical comparison results are provided in the supplementary materials.
6.2. Participants and Apparatus
@@
We used a Meta Quest 2 headset and stock controller, reflecting a typical consumer-grade VR setup, powered by a desktop featuring an NVIDIA RTX 3080 GPU and an AMD Ryzen 5 5600G CPU. This setup provided consistent performance across all trials, minimizing hardware-induced variability.
6.3. E1: Pick-and-place — Method
@@
Figure 5. RQ scores for the Pick-and-Place task, using a 7-point Likert scale ranging from -3 (strongly disagree) to +3 (strongly agree). Brackets indicate statistical significances (* : $p$ $<$ .05, **∗ ∗ : $p$ $<$ .01).
@@
Box plots showing RQ scores for the Pick-and-Place task across four categories: Total Score, Motor Control, Finger Movement Realism, and Interaction Realism. Each category compares ForceGrip, VR-HandNet, and Attachment methods. ForceGrip consistently shows higher median scores. Statistical significance between methods is indicated with asterisks.
6.4. E1: Pick-and-Place — Results
[…]
Participants provided qualitative observations that illuminated each technique’s strengths and limitations. For FG, participants praised it’s ability to “stably and realistically grasp small objects” (P3, P7, P12) and “naturally adapt finger motions to varied shapes” (P6, P9, P11, P13, P14, P19, P20). Some, however, experienced occasional “object slipping” (P3, P6, P10), likely due to the overly rounded fingertip shape of the capsule collider model provided by the SDK.
@@ 6.4. E1: Pick-and-Place — Results
Overall, this experiment indicates that ForceGrip provides superior visual realism for typical object manipulation tasks. Examples of user interactions with various object sizes and shapes, supporting these findings, are shown in Figure 10 and Figure 11.
6.5. E2: Can squeeze — Method
@@
Participants interacted with a virtual can modeled after a standard beverage can and were asked to deform it to a target level from 0 (none) to 10 (maximum). Each discrete level represented a 1 kgf integer increase. We used this 11-level scale to balance force granularity with the limited sensitivity of VR triggers, as finer subdivisions made it difficult for participants to modulate shallow trigger inputs in preliminary tests.
At the start of each trial, a can appeared in front of the user along with a random target deformation level on the display. The participant adjusted the VR controller’s trigger to deform the can to match the target level, then placed the can into a container. Visual and auditory feedback indicated deformation through mesh updates and sound effects.
@@
Figure 6. FRQ scores for the Can Squeeze task, using a 7-point Likert scale ranging from -3 (strongly disagree) to +3 (strongly agree). Brackets indicate statistical significances (* : $p$ $<$ .05, **∗ ∗ : $p$ $<$ .01).
[…]
Box plots showing FRQ scores for ForceGrip and VR-HandNet across four metrics: Total Score, Motor Control, Grip Force Realism, and Interaction Realism. ForceGrip shows consistently higher median scores with narrower variability. Statistically significant differences are marked with asterisks.
6.6. E2: Can squeeze — Results
As shown in Figure 6, there exist significant differences between FG and VH across all metrics. There was no statistically significant change in SSQ scores before and after the experiment.
@@ 6.6. E2: Can squeeze — Results
Figure 7. Correlation between trigger input ($x$) and grip force ($y$) for FG and VH, with the reference line ($y=10x$). Outliers (top 5%) were excluded, and the remaining data was duplicated and resampled to 1,000 points, ensuring an even distribution of trigger values. For FG, points below the reference line mostly occurred when the object slipped from the agent’s grasp.
@@
To visualize how closely each system reflects user intentions, we plotted the trigger input against the measured grip force (Figure 7). FG exhibited a clear linear correlation ($r=0.884$). In contrast, VH showed negligible correlation ($r=0.073$), consistently producing excessive grip forces irrespective of trigger amplitude. This result contradicts the 1.2 kgf maximum force claimed by (Han et al., 2023) and supports reports that VH appears to lack grip force control.
7. Conclusion and Future Improvements
@@
We have introduced ForceGrip, a novel approach for delivering realistic hand manipulation in VR controller interfaces through faithful grip force control. Unlike many prior systems that rely on reference datasets, ForceGrip leverages automatically generated training scenarios. Our curriculum learning framework further addresses the challenges posed by complex, dynamic interactions, enabling robust convergence and state-of-the-art performance across diverse virtual manipulation tasks.
One significant limitation of our method is its reliance on a one-dimensional control signal in the range of [0, 1], chosen for broad compatibility. While this simplifies user input, it limits complex actions such as in-hand manipulation or tool-specific adjustments (e.g., scissors, chopsticks), which would require a richer input structure. Improving collision and friction modeling through soft-body simulations could produce more stable and lifelike contact dynamics, particularly for slippery or deformable objects.
Acknowledgements.
This work was supported by the Sponsor National Research Foundation of Korea https://www.nrf.re.kr/eng/main under Grant No. Grant #RS-2025-00518643 (70%) and Grant #RS-2025-00564137 (10%). It was also supported by the Sponsor Institute of Information & Communications Technology Planning & Evaluation (IITP) https://www.iitp.kr/en/main.it grant funded by the Korea government (MSIT) under Grant No. Grant #IITP-2025-RS-2020-II201819 (10%) and Grant #RS-2020-II200861 (10%).